To determine the number of deer in a game preserve, a conservationist catches 573 deer, tags them and lets them loose. Later, 840 deer are caught; 210 of them are tagged. How many deer are in the preserve?
I understand that you need to subtract the 210 from the 840, since they were already tagged...would you also subtract them from the 573 that were previously tagged and then add all the newly tagged ones to that answer?
Feel free to email me at shannon_corbin@live.com

Answer by jim_thompson5910(35256)

(Number of deer Total)/(Number Tagged) = (Number of deer Total)/(Number Tagged)

x/573 = 840/210

210x = 573*840

210x = 481320

x = 481320/210

x = 2292

So there are 2292 deer in the preserve.